Projection Admin 2/5 Lillian (Alec on mid)



To do this evening/ was completed: There is a headset in T5 at seat 306. Please pick this up once the feature is over. Thanks!

Presentation Issues:

Preshow Clips/Playlist: Took a look at The Moment today, which I brought up to 5.5 as some dialogue can be soft, but the music is loud enough. 

We're moving the Superbowl to T2, swapping the 6:45 The Moment and 9:15 Send Help shows to 3. Made this adjustment on the schedule and cue sheet for Sunday. -A

Trailer Updates:

Special Events/ AV set up: Setup for Spoons today in T1. Tested the wireless mics, laptop and spots. All you need to do is turn everything on, plus the breaker for the spots and bring down the mics from their charging ports once it is time. The mic stands are by the table, also set up.In addition to this, the three mics and chairs are all set for Removal of the Eye in T3, which I got eyes on just to double check.
--Made sure they were all set before heading out. Tested mics, picture, sound, house and spot lights. All good. -A

Ordering/Receiving (posters & trailers): Super Mario posters stored and labeled. 

Ordering/Receiving (content): Currently unzipping the DCP of Make Out With Violence, which should then get put in the Share folder. For future reference, try to download DCPs directly to the Share folder. Thanks! 
--I restarted the extraction with Keka, it seemed like the Unarchiver was really slowing the computer down. Also, Keka gives you the option to extract to a new location, so it unzipped right into the Share folder. -A 
Dead Lover successfully ingested to T3 and got eyes on this - everything looks/sounds good. Gave Joe some downloading suggestions since he's been having a hard time with it at WB but I think he's gonna use the one Kat made. 
Ingested Goat and V8 all ages preshow to the LMS. Ingested Corpse Bride to the LMS. 
Coach Carter slide DCP'd and ingested. 
We received Coach Carter and A Time to Kill today. Both in T1. 

Ingested Make Out with Violence to the LMS -A

Outgoing/Shipping (content): Scheduled FedEx express pickup for Unbreakable and Die Hard, which I handed off to the FedEx person. These are now outbound. Tracking is in the asset sheet. 
Got rid of playlists/DCPs we do not need. 

Had our last round of Is This Thing On? today, which I have since deleted from T6. 

Return label printed and attached to the Let the Right One In print, on the cart. Pickup is scheduled for tomorrow. -A

Print/video Inspection: 

Asset tracking doc updated/checked?: Yes

Maintenance/cleaning: Alec helped me sort out the last reel for Let the Right One In (thank you, Alec) which is now safely packed up with our own shipping reel. 

Cue sheets updated/checked (every shift): Yes. Info for coming days updated today. 

Marquee updated?: Added The Moment to the marquee, Alec did the poster. 

Posters updated?: Is This Thing On? poster and letters are down, Pillion poster is now outside with a coming soon banner. -A

Cleaning lights on for the night?:

Projection cart by entrance?:

Additional Notes:

Comments

Popular posts from this blog

NPP Projection Admin (Lillian/Will) 7/7

Projection Admin 9/8 Lillian/Will

NPP Projection Admin 7/6